OpenStudy (anonymous):

A baseball is thrown straight up from the rooftop 192 feet high. The function s(t) = -16t2 + -64t + 192 describes the ball's height above the ground, s(t), in feet, t seconds after it was thrown. How long will it take for the ball to hit the ground?

OpenStudy (anonymous):

Are we to assume that the ball is thrown slightly over the edge of the roof?

OpenStudy (anonymous):

this is all the question said, so i'm not sure :( can u still try to help?

OpenStudy (anonymous):

I'm going to assume yes. This gives us a quadratic equation: 0=-16t^2-64t+192 a.k.a. 0 = t^2+4t-12

OpenStudy (anonymous):

The negative solution we can ignore, the positive is your answer.

OpenStudy (anonymous):

(t+6)(t-2)=0 Therefore t=2

OpenStudy (anonymous):

ok.. thank you